The students run around a rectangular court that is 3 m wide and 4m long. How far did the students run?

Answer by jim_thompson5910(28715)

Perimeter = 2*(Length + Width)

P = 2*(L + W)

P = 2*(4+3)

P = 2*7

P = 14

So they ran 14 meters